### v1

# JWST Reveals Varied Origins Between Jupiter’s Irregular Satellites

Abstract
We report observations of eight Jovian irregular satellites with JWST’s NIRSpec instrument: Himalia, Elara, Pasiphae, Sinope, Lysithea, Carme, Ananke, and Themisto. Irregular satellite families, which are presumed to have formed via collisions, contain various Trojan-like and C-type-asteroid-like surfaces. We sample the three largest members of the Himalia satellite family, detecting the presence of complexed CO2 and a unique absorption band from whose character correlates with satellite size. The two largest irregular satellites, Himalia family members Himalia and Elara, contain ammoniated phyllosilicates that are not seen in the meteorite inventory. We propose that the Himalia parent body was heterogeneous and formed with materials similar to Ceres-like ammonium-bearing asteroids. Several small () irregular satellites closely track the colors and absorption bands of “red” Jovian Trojans, demonstrating that these compositions are retained amongst the products of collisions that occurred after Jovian capture. We report the first detection of aqueous alteration products in the retrograde satellite swarm, finding Ananke’s 3 band to closely match phyllosilicates seen in C2 chondrites. Notably, objects with OH absorption features similar to the Trojan asteroid Eurybates are found in both the retrograde Pasiphae family and the prograde Himalia family, confounding a simple link between such materials and a single surface type. The irregular satellites appear consistent with some materials that experienced alteration from liquid water and others that did not. Consequently, Jupiter may have captured bodies that formed from different initial compositions, or bodies that experienced different levels of heating, driving differential alteration processes.
1 Jupiter’s Irregular Satellites: Possible Remnants of Outer Solar System Mixing
The origins of Jupiter’s irregular satellites are uncertain. The highly inclined and eccentric orbits of these small satellites lie exterior to the regular, Galilean satellites, suggesting that they were captured from asteroid- or comet-like parent reservoirs (e.g., Sheppard & Jewitt, 2003; Nesvorný et al., 2007, 2014). Jovian irregular satellite orbits are also clustered, with one large prograde group (Himalia), three major retrograde groups (Ananke, Carme, and Pasiphae), and several smaller, isolated satellites (Sheppard & Jewitt, 2003). The (a,i) orbital elements of the Jovian irregular satellites, noting major family clusters, are displayed in Figure 1.
Various processes have been proposed to explain the capture of Jupiter’s irregular satellites, as well as similar populations at Saturn, Uranus, and Neptune. Nesvorný (2018), Jewitt & Haghighipour (2007), and Nicholson et al. (2008) provide thorough reviews of these mechanisms. Some relate the capture process to the formation of the planet itself, including pull-down capture (Heppenheimer & Porco, 1977, where the planet’s mass growth captures nearby resonant objects) and gas-drag (Pollack et al., 1979; Astakhov et al., 2003; Ćuk & Burns, 2004; Kortenkamp, 2005, where nebular gas affects nearby planetesimals that drift into bound states). Alternatively, three-body interactions could emplace satellites after planet formation (Colombo & Franklin, 1971).
Notably, models of planet migration reproduce the overall number and orbital distributions of each giant planet’s irregular satellites via three-body interactions between a (to be captured) small body and two closely interacting planets (Nesvorný et al., 2007, 2014). Irregular satellite capture may therefore relate to large-scale migratory processes that have sculpted the overall architecture of the solar system. Similar capture models can explain the present-day Jupiter Trojans (Nesvorný et al., 2013), and planet migration provides key insight to understand the main asteroid belt (e.g., Minton & Malhotra, 2009; Morbidelli et al., 2010; Walsh et al., 2011), Pluto’s orbit (Malhotra, 1993), and the complex populations in the trans-Neptunian region (see recent review by Gladman & Volk, 2021).
Through the lens of migratory capture, the most likely source reservoir for irregular satellites and Trojans would be the primordial Kuiper belt. This is a hypothesized region stretching from and au (Nesvorný et al., 2019), immediately outside the locations where the giant planets accreted ( and au, e.g., Tsiganis et al., 2005; Nesvorný & Morbidelli, 2012; Nesvorný et al., 2019). Within Myr after the dissipation of the solar nebula, Neptune’s migration through the primordial Kuiper belt would have destabilized this planetesimal population and triggered dynamic instability throughout the outer solar system (Tsiganis et al., 2005). Planetesimals ejected inwards would enter planet-crossing orbits, and those ejected outwards would create the scattered disk (Levison et al., 2008). Meanwhile, the present-day dynamically ”cold” Kuiper belt population would represent the portion of the primordial Kuiper belt that formed far enough away from the giant planets to remain stable (although this population may have also been largely depleted in mass, see e.g. Gomes et al., 2018)
While the irregular satellites could have been captured from a single source, they could also represent a composite of various populations that interacted with Jupiter over the planet’s history. The three-body capture scenarios defined by Nesvorný et al. (2007) occur in the vicinity of primordial Kuiper belt remnants, not the original generation of middle solar system planetesimals. However, some closely-bound irregular satellites may be able to survive from previous capture events. For example, Nesvorný et al. (2014) find that of Himalia-like Jovian satellites could survive planetary encounters that occurred after their capture. The likelihood of such a scenario is currently speculative, but can be constrained by determining the variety of the present-day irregular satellite population. Determining whether or not all irregular satellite parent bodies formed within similar primordial environments provides a key constraint on the character of Jupiter’s migratory history.
Observational evidence does not yet uniquely discriminate irregular satellites’ formation region(s). Jupiter’s position, proximate to both the inner solar system’s reservoirs of asteroids as well as the outer solar system’s comets, trans-Neptunian objects (TNOs), and Centaurs, provides a rich array of comparison populations. Determining the irregular satellites’ compositions in the context of these other populations tests how migratory capture could have occurred. But the post-capture evolution of the irregular satellites, including surface alteration and collisional breakup, must be accounted for to forge successful cross-population links. To determine the lineage(s) of Jupiter’s captured satellites, we therefore require knowledge of their bulk compositions and the variability within various collisional families. We provide these bulk compositional assessments with high-SNR JWST spectra of the largest objects in each each major satellite family, and we provide the first detailed exploration of variability within a family as a function of size. These constraints address irregular satellite formation conditions and the evolutionary processes that have shaped them.
1.1 Previous Observational and Laboratory Context
Numerous spacecraft missions have explored the Jovian system, but only the largest irregular satellite, Himalia, has been the subject of resolved spacecraft imaging (Porco et al., 2003, resolved across 5 pixels) or spectroscopy (Brown et al., 2003, SNR). Telescopic observations have been the primary method to understand the population. These characterization efforts suggest kinship between irregular satellites and various groups of dark primitive asteroids. Brown & Rhoden (2014) found that Himalia has a spectral feature that indicates the presence of hydrated material and appears similar to the main belt asteroid (52) Europa (Takir & Emery, 2012). This feature is only partially observable from ground-based telescopes, which limits compositional tests beyond that comparison. Himalia and its associated family were also reported to display a 0.7 feature of variable strength correlated with size (Vilas & Hendrix, 2024).
Amongst other satellite families, numerous neutral-to-very-red spectra have been observed, similar in color to Jupiter Trojan asteroids and without apparent absorption features at wavelengths shorter than (Sharkey et al., 2023; Bhatt et al., 2017). Coarser color photometry surveys (Graykowski & Jewitt, 2018; Grav et al., 2003; Grav & Holman, 2004) have characterized a wider satellite sample and have found Jupiter’s irregular satellites to have colors similar to those of Saturn, Uranus, and Neptune. But the relative scarcity and faintness of Jupiter’s irregular satellites have limited further attempts to characterize them. These observational constraints generally limit even photometric studies to the few Jovian satellites that are larger than D 10 km, and they have precluded the high-SNR spectroscopy that is needed to interpret mineral absorption features.
Existing high-resolution, high-SNR spectra of Galilean satellites may provide an alternate means to investigate irregular satellite compositions. Collisions within irregular satellite swarms generates dust that likely collides with the surfaces of the Galilean moons, in particular the leading hemisphere of Callisto (e.g., Bottke et al., 2010, 2013, 2024; Chen et al., 2024). Accumulated dust from these primordial bodies then mixes with native components in the Galilean moons’ regoliths that is subsequently irradiated by charged particles trapped in Jupiter’s massive magnetosphere (Johnson et al., 2004). This process would form new components, such as CO2 (e.g., Gomis & Strazzulla, 2005), sulfates (e.g., Carlson et al., 2002), and perhaps complex CN-bearing organics (e.g., Cartwright et al., 2024) depending on the nature of the exogenic material.
However, the lack of high quality spectral datasets for Jupiter’s irregular satellites limits the scope of hypothesis testing both for irregular satellite compositions and their connections to the Galilean moons. In particular, a broad band centered near 4.57 on Callisto’s leading hemisphere has been speculated to have an irregular satellite origin (Cartwright et al., 2020, 2024), but this speculation cannot be tested without searching for the same feature on irregular satellites. Without identifying specific classes of silicate minerals, ices, or organic species, the origins of the irregular satellites, and their effects on the Jovian system, remain mysterious.
Laboratory studies of meteorites further reinforce the need to characterize absorption features to properly identify minerals, ices, or organics on irregular satellites. Petrologic, isotopic, and mineralogical study of the meteorite Tarda suggests that it is a sample from a parent body that is either the same or closely related to that of the Tagish Lake meteorite. This link is particularly interesting given that Tarda’s spectrum is “less-red” while Tagish Lake is “red” (Schrader et al., 2024). More generally, a variety of carbonaceous chondrites with wildly different alteration histories and compositions (petrographic grades 1-3) can have remarkably similar colors from 0.7-2.5 (e.g., Cloutis et al., 2011a, b). Strikingly, the color variation within individual sample preparations of some CM and CI chondrite samples can exceed the differences between C/P/D-type asteroids (see packing conditions of various spectra of CI Alais in Cloutis et al. 2011a; see grain-size variations in both CM2 Aguas Zarcas and Tarda from Cantillo et al. 2021 and CM2 Murchison by Cloutis et al. 2018). Given the magnitude of these noncompositional spectral effects, Sharkey et al. (2023) noted that differences in the activity histories between red/less-red objects (speculatively caused by different volatile abundances) could be a confounding driver of their color differences, even if their present-day surfaces contain similar materials.
With the advent of JWST, our understanding of small solar system bodies is rapidly expanding. JWST’s NIRSpec instrument’s high sensitivity over key infrared passbands identifies assemblages of rocky minerals, volatile ices, and organic compounds with spectroscopic complexity previously accessible only via laboratory study. Recent observations of Jovian Trojans by Wong et al. (2024) highlight the importance of studying carbonaceous asteroids from 0.7-5.2 . When observed from Earth-based telescopes, Trojans are featureless 2.5 . When observed via JWST, distinct spectral features from 2.5-5.0 in both color groups were discovered, providing the first identifications of organic compounds and OH-bearing materials on these asteroids.
The Wong et al. (2024) characterization of 3 bands on Trojans provides new insight into the evolution and origins of red/less-red surfaces. If Jovian Trojans formed originally in the primordial Kuiper belt, they would have lost any surface ices that become volatile between 5-40 au during their inward migration. But more complex processing could have also occurred to produce stable compounds from volatile mixtures, such as hypothesized space weathering effects (e.g., Wong & Brown, 2016). The irradiation of organic and ice-rich materials have been demonstrated to create complex assemblages in laboratory simulations of asteroidal surfaces (Urso et al., 2020). Such scenarios have been hypothesized to explain the rounded 3 absorption features on some main-belt asteroids, suggesting that they may have formed in outer solar system reservoirs under conditions cold enough to incorporate methanol ice and ammonium (Rivkin et al., 2022). This hypothesis may also apply to main-belt comets like 238P/Read, which has a similarly rounded band (Kelley et al., 2023), as well as Jupiter Trojan asteroids with similarly shaped absorptions (Wong et al., 2024).
Coupled with their shared heliocentric distance, the strong spectral slope similarities between the Jovian Trojans and irregular satellites already suggest some level of compositional similarity. However, the effects of more frequent collisions within the irregular satellites (Bottke et al., 2010), and the presence of objects similar to outer-main belt asteroids (Brown & Rhoden, 2014; Takir & Emery, 2012) complicates a direct equivalence between the two populations. This work details the mineralogies of Jupiter’s irregular satellite groups, shedding new light on the organic-rich material reservoirs that have interacted with Jupiter and the Galilean moons.
2 Observations and Data Reduction
Satellite | Diameter (km) | Observation Date | Phase Angle | Sun-Target | Target-Observer |
(Degrees) | Distance (au) | Distance (au) | Distance (au) | ||
JXVIII Themisto | 8aaVia Sheppard & Jewitt (2003), assumes geometric albedo of 0.04. | 2024-01-31 | 11.59 | 4.94 | 4.90 |
JVI Himalia | bbVia (Grav et al., 2015) from NEOWise thermal-IR observations. | 2024-01-31 | 11.54 | 4.96 | 4.92 |
JX Lysithea | bbVia (Grav et al., 2015) from NEOWise thermal-IR observations. | 2024-01-25 | 11.55 | 4.96 | 4.82 |
JVII Elara | bbVia (Grav et al., 2015) from NEOWise thermal-IR observations. | 2024-01-18 | 11.40 | 4.97 | 4.72 |
JXII Ananke | bbVia (Grav et al., 2015) from NEOWise thermal-IR observations. | 2024-01-31 | 11.19 | 5.11 | 5.08 |
JXI Carme | bbVia (Grav et al., 2015) from NEOWise thermal-IR observations. | 2024-01-25 | 11.92 | 4.81 | 4.69 |
JVIII Pasiphae | bbVia (Grav et al., 2015) from NEOWise thermal-IR observations. | 2024-01-25 | 11.53 | 4.97 | 4.89 |
JIX Sinope | bbVia (Grav et al., 2015) from NEOWise thermal-IR observations. | 2024-01-21 | 11.58 | 4.94 | 4.76 |
Observations were performed using the NIRSpec integral field unit (IFU) in prism mode () for all targets except Himalia, which was observed in the higher resolution G235M and G395M grating modes (). Observational circumstances, including timing and geometry, are given in Table 1.
Calibration and spectral extraction of the Jovian irregular satellite data proceeded similarly to the processes described in Souza-Feliciano et al. (2024); De Prá et al. (2024); Licandro et al. (2024); Pinilla-Alonso et al. (2024). This process is similar to that described in Wong et al. (2024), with some differences. To start, uncal files for all targets were downloaded from the Barbara A. Mikulski Archive for Space Telescopes (MAST) using the jwst_mast_query tool and processed through the JWST data calibration pipeline locally. Version 1.14.0 of the calibration pipeline (Bushouse et al., 2024) and reference context jwst_1225.pmap were used. The pipeline was run with default parameters and included the non-default NSClean step to remove the 1/f pattern noise (Rauscher, 2024). Processing was performed up to Level 2b and the production of dither-specific s3d spectral data cubes.
Spectral extraction on each individual data cube made use of the “template PSF-fitting” routine first described in Wong et al. (2024). The uniformly spaced wavelength array was constructed using the CRVAL3 and CDELT3 header keywords and slices 0.7 m and 5.1 were trimmed due to higher noise. Then, the DQ extension was used to identify pixels flagged by the pipeline and any pixel with a non-zero value was set to NaN for the remainder of the extraction. Target centroids were first determined by-eye (typically good to a few tenths of a pixel) and used to measure radial distances from the target. Outlier pixels due to artifacts or background objects, as well as those 5 pixels from the centroid were set to NaN. Then the background was calculated in each slice of the cube using the median of all pixels 5 pixels from the centroid and subtracted from that slice. A template PSF was then computed as the median of the 10 background-subtracted slices proceeding and succeeding the slice under consideration (inclusive) for a total of 21 slices. Slices near the edge of the considered wavelength range included 21 total slices, but the resulting template PSFs were still adequate. The template PSF was then trimmed to a 99-pixel vignette, with all external pixels set to NaN, in order to reduce the leverage of more distant pixels when fitting to the central slice in the set. The integrated flux of the vignette was then normalized to unity. The scipy.optimize.minimize function and a Nelder-Mead (amoeba) algorithm was used to determine the best-fit values for the flux scaling factor of the template PSF and the background level. This fit was iteratively performed 3 times, using the outputs from the previous iteration as the initial guess for the next iteration. Additionally, after each iteration a model was constructed (template PSF multiplied by the flux scaling factor followed by the addition of the background) and subtracted from the data, with any pixels 5- from the mean counted as outliers and set to NaN prior to the next iteration. After the third iteration, a center-of-mass centroid was computed to determine the target centroid to sub-pixel accuracy, the template PSF was multiplied by the flux scaling factor, and the flux was extracted within a 3.5-pixel radius circular aperture.
Following the construction of the 1D spectra as described above, the individual dithers were combined and the reflected solar component was divided out. The G-type standard star SNAP-2 (Gordon et al. 2022), observed in program 1128 (PI: Lutzgendorf), was used for this purpose and these data cubes underwent the same processing as described above for the irregular satellites data cubes. All individual standard star spectra were first resampled onto the same wavelength grid and normalized by the median of all dithers. The purpose of this resampling and normalization was to reduce the spurious spread in absolute flux values due to placement of the target at different locations within the IFU field of view. The median was then computed in each wavelength bin and multiplied by the previously calculated median to retrieve the absolute flux scaling. Three iterations of outlier rejection were performed with a 21-point moving median and a 3- threshold; any points found to exceed this threshold were replaced with the value of the moving median. The irregular satellite 1D spectra underwent the same resampling, normalization, median, rescaling, and outlier rejection steps and then were divided by the SNAP-2 spectrum. Uncertainties on each medianed spectrum were computed as the median absolute deviation of each wavelength bin and uncertainties were propagated following division by the standard star.
Our processed, flux calibrated spectra include a component of thermal emission that begins near 4.0 microns and increases in strength to dominate observed fluxes 4.5 microns. The thermal emission component manifests as a broad, blackbody-like curve that distorts our analysis of reflected light. To isolate the reflected component of the observed flux, we estimated the thermal component via an iterative implementation of the near-Earth Asteroid Thermal Model (NEATM, Harris, 1998; Delbó & Harris, 2002). This model estimates thermal flux from an asteroidal surface by accounting for observing geometry, asteroid size (diameter), albedo, and a beaming parameter, , which broadly accounts for variations according to different surface-scattering scenarios. For the purposes of our spectral corrections, we took the geometric albedo and as free parameters and took values of satellite diameters from Grav et al. (2015).
The NEATM models were fit using emcee, a Markov-Chain Monte Carlo package in python (Foreman-Mackey et al., 2013). We took albedo values of 2% and values of 1.0 as our initial solutions for all objects. Our goals for these models are not to provide detailed explorations of thermal model parameters (including their possible degeneracies), but instead to estimate the shape of the thermal flux curve. To do so, some assumption about the shape of the underlying reflectance spectra is required. We estimated the reflected component of each satellite spectrum as a linear extension of the observed spectral slope at the cut-on point of the thermal emission (variable, near 4 microns). MCMC fits were then performed to recover NEATM model solutions and best-fit parameters. Specifically, fits were iterated to ensure good agreement to the cut-on wavelengths and the magnitudes of peak thermal emission. The choice of baseline was iterated until a smooth output spectrum was produced, to ensure that spectral slopes were not extrapolated from regions with absorption bands. Figure 2 shows a representative example of a finalized thermal flux estimate and the NEATM best-fit model for Lysithea. Reflectance spectra were then computed by subtracting the NEATM best-fits from the original flux-calibrated spectra.
3 Spectral Groupings and Band Identifications
3.1 Continua Definitions
The processed, thermally-corrected reflectance spectrum of each target is shown in Figure 3 (Left). Proper analysis of absorption features requires clear definitions of the spectral continuum. To aid reader comparisons, we define a “global” continuum, used in constructing Figure 3 (Right). This continuum was fit per-object with a low order (2-4) polynomial on either side of the identified 3 micron complex. Dividing the spectrum by this polynomial transforms the spectra from relative reflectance to continuum-divided reflectances. In continuum-divided space, spectral features are identified via strong divergence from unity.
This global continuum is well-suited to identify major features, but it is imprecise for characterizing weak and/or narrow absorption bands. We report band centers and depths in Table 2 by defining local continua as follows. For the broad 3 band, a linear continuum was fit from 2.7 to 3.8 microns, with the centers then identified by a polynomial (3rd degree) fit to the bottom half of the band. The listed depth is from this linear continuum to the center position. We additionally report the band depth at 2.9 to facilitate comparisons with ground-based observations.
All other bands are quantified by fitting a local linear continuum on either side of the feature, then fitting a Gaussian to identify the center and depth from the local continuum. Due to its asymmetry, we quantify Himalia’s 2.7 feature uniquely. Its center is defined by fitting a Gaussian near its sharp minimum, and its depth is defined relative to the linear continuum of the overall 3 band complex.
To determine the robustness of individual features, band parameters are reported according to a Monte Carlo resampling procedure. After continuum division, individual data points are resampled according to their Gaussian uncertainties before fitting feature depths and centers. Quoted parameters are the mean of 10,000 resampling iterations, and the standard deviations of the resampled parameters taken as the uncertainty.
3.2 Spectral Groups
We find that the irregular satellites display a variety of spectral characteristics, with some groupings clustered within individual families and some that appear in multiple families. We describe four spectral categories to illustrate these complex relationships. Band parameters for all features observed on individual satellites are given in Table 2. The first group is represented by Sinope, Carme, and Themisto, which are all Trojan-like objects all in different orbital families. Ananke comprises the second group, displaying a deeper, rounded 3.0 micron band. The third group includes Pasiphae, a retrograde satellite, and Lysithea, a prograde satellite, which have similar 3.0 micron features with several weak local minima between 2.7-3.3 . Finally, Himalia and Elara, both in the prograde Himalia family, show a unique band complex that includes a sharp 2.7 micron band transitioning into a broad, round feature from 2.8-3.5 microns.
Sinope and Carme have similarly steeply increasing spectra from 0.7-2.5 microns, and shallow absorption features near 3.0 and 3.4 , akin to the “red” class of Jupiter Trojans (Emery et al., 2011; Wong et al., 2024). Notably, Themisto, a small () individual prograde satellite interior to the Himalia family, groups closely with these more weakly bound retrograde satellites. The spectral slopes of these objects, when overlapping, are consistent with previous measurements at VNIR wavelengths by Sharkey et al. (2023). Carme has a slightly deeper band at 3.0 microns than Sinope and Themisto (depths of vs. ), but they all have similar 3.4 depths ().
Carme | Sinope | Themisto | Ananke | Pasiphae | Lysithea | Elara | Himalia | ||
Center () | - | - | - | - | |||||
Band | Depth (%) | - | - | - | - | ||||
Attribution | - | - | - | ? | - | Carbonates? | ? | ? | |
Center () | - | - | - | ||||||
Depth (%) | - | - | - | aaHimalia band depth is from the top continuum across the 3 band, not local. All others are local. | |||||
Sharp Band | Attribution | - | - | - | ? | ? | ? | Mg-rich | Mg-rich |
Phyllosilicate | Phyllosilicate | ||||||||
Center () | |||||||||
Band | Depth (%) | ||||||||
2.9 Depth | |||||||||
Center () | bbSinope organics band may have two components – and ). Values in the table assume a single band. | - | - | ||||||
Band | Depth (%) | - | - | ||||||
Attribution | Organics | Organics | Organics | Organics | Organics | - | Organics | - | |
(Aliphatic) | (Aliphatic) | (Aliphatic) | (Aromatic) | (Aliphatic) | - | (Aliphatic) | - | ||
Center () | - | - | - | - | - | ||||
Band | Depth (%) | - | - | - | - | - | |||
Attribution | - | - | - | - | - | ? | Carbonates? | ? | |
Center () | - | - | - | - | - | - | |||
Band | Depth (%) | - | - | - | - | - | - | ||
Attribution | - | - | - | - | - | - | Complexed | Complexed | |
Center () | - | - | - | - | - | ||||
Band | Depth (%) | - | - | - | - | - | |||
Attribution | ? | ? | ? | - | - | - | - | - |
We also identify possible absorption features at 4.9 on Carme, Sinope, and Themisto. These features are within lower-SNR spectral regions, which are affected by thermal tail removal that impacts the precision of continuum definitions. If real, this feature is similar to one observed by JWST on the surface of Callisto (Cartwright et al., 2024). Its origin is speculative, but could relate to radiolytic production of from , or the C-O symmetric stretching mode () of carbonyl sulfide (OCS). OCS is a common sulfur bearing species observed in interestellar ices in dense molecular clouds (Palumbo et al., 1997; Ferrante et al., 2008; McClure et al., 2023) and cometary comae (Saki et al., 2020; Biver & Bockelée-Morvan, 2019).
Ananke is unique within our sample with a deep, rounded 3.0 complex and the most neutral (least-red) slope we observe. Its 3 band depth () is similar to the redder Pasiphae and Lysithea, though their band centers and shapes differ. A narrow feature at 2.86 is overprinted within the rounded 3 complex. Beyond 3.6 , Ananke’s spectrum is neutral and featureless. Our observations of Ananke differ from the ground-based observations of Sharkey et al. (2023) from 0.7-1.6 microns, which showed a slope that is approximately 10% redder. A similar discrepancy was noted by Wong et al. (2024) when comparing their JWST observations of the Trojan Polymele to Sharkey et al. (2019)’s ground-based observations with the same methodology and setup. Given that both Polymele and Ananke have similar brightnesses V 19, it is possible that those short wavelength slope discrepencies are indicative of systematic errors in the ground-based observations of these very faint objects.
Pasiphae (Pasiphae family) groups closely with Lysithea (Himalia family) in terms of spectral slope and absorption features. They trace closely together shortward of . Pasiphae displays a steeper red slope from 3.5 to 5.2 , while Lysithea rolls over to neutral (flat) reflectance, similar to Ananke. Pasiphae and Lysithea display deep absorption features () near 3.0 that are distinct from any other satellites.
A comparison of Pasiphae’s and Lysithea’s continuum-divided bands are shown in figure 4, highlighting the similar absorption shapes between them. Both Pasiphae and Lysithea have additional faint features of 3% overprinted near 3.01 . Lysithea has an additional feature near 3.9 , which is similar to features caused by carbonates. However, carbonates also display absorptions near 3.4 which we do not detect, confusing a clear identification. Lysithea also shows a narrow feature at 2.63 whose origin is less clear, but is consistent with features seen in some carbonates (Bishop et al., 2021).
Himalia and Elara, both in the prograde Himalia family, are largely similar except for their 2.7 bands. They have similar spectral slopes from 1.6-5.2 , and similarly located major absorption features at 2.7 and 3.05 . However, the relative strength of those major absorption features differs strongly, with Himalia showing a much deeper 2.7 minimum () than Elara (). Our present spectrum of Elara from 0.7-1.6 also compares closely with ground-based spectra of Himalia from Bhatt et al. (2017). The band complexes in these objects indicate the presence of phyllosilicates and are discussed further in Section 4.
We also find a feature at 4.25 on Himalia and Elara. This feature is similar to one discovered on the Jupiter Trojan Eurybates by Wong et al. (2024). Figure 5 compares our observations of Himalia and Elara with the the Wong et al. (2024) observations of Eurybates, as processed via our reduction pipeline. We reprocess this spectrum only for methodological consistency between these datasets but do not note any differences between reductions. This feature is similar to those caused by solid-state CO2, which is noteworthy since CO2 is volatile and unlikely to be stable as a crystalline ice on the relatively warm airless surfaces of Jovian satellites or Trojans (Wong et al., 2024). Since CO2 is implausible as pure surface ice, it is likely complexed with more refractory components, similar to the Galilean moons (e.g., McCord et al., 1998; Villanueva et al., 2023; Cartwright et al., 2024). The discoveries of the 4.25 feature on Eurybates, Himalia, and Elara demonstrates that complexed CO2 is present on a variety of objects captured by Jupiter.
The 3.0 absorber(s) for Carme, Sinope, Themisto, Pasiphae, and Lysithea are consistent with a variety of OH-bearing species, which we discuss in Section 8 in the context of meteorite comparisons. The absorption at 3.4 on Elara, Pasiphae, Ananke, Themisto, Sinope, and Carme is consistent with complex organics. All three objects in the Himalia family (Himalia, Elara, Lysithea) display a feature near 3.9 . On Elara, the 3.4 and 3.9 features may be consistent with carbonates, but with an unclear relationship to the 3.9 absorbers on Himalia and Lysithea.
4 Ammoniated Phyllosilicates in the Himalia Family
Himalia’s complex spectrum from 2.6-3.8 strongly constrains its phyllosilicate composition. The sharp and deep feature at 2.7 can be reproduced with a variety of hydrated silicate materials from terrestrial and meteoritic inventories (e.g., Takir et al., 2013, 2019; De Angelis et al., 2021). However, none of the hydrated silicates that match Himalia’s 2.7 feature have the additional rounded minimum at 3.05 microns.
Some meteorite materials with significant amounts of adsorbed terrestrial water can reproduce the overall width of Himalia’s band complex. However, we note that experimental studies of both meteorites and terrestrial phyllosilicates find that adsorbed water, and the spectral feature it causes, is easily removed when samples are treated under proper vacuum and desiccation procedures designed to mimic asteroid surface conditions (Beck et al., 2010; Takir et al., 2013, 2019). We therefore disfavor fits to Himalia that use phyllosilicates containing adsorbed water.
Himalia’s combination of absorptions at 2.7 and 3.05 is reminiscent of Ceres (Figure 6), where the 3.05 features is associated with the presence of NH4 in the clay mineral structure (e.g., King et al., 1992; Rivkin et al., 2006; Ammannito et al., 2016; de Sanctis et al., 2016; De Sanctis et al., 2017). Both Himalia and Ceres have 2.7 bands with consistent centers (2.72 ; de Sanctis et al., 2015). Pilorget et al. (2021) also detected this combination of 2.72 and 3.05 bands within the sample of asteroid Ryugu returned by Hayabusa2 mission. The Ryugu sample also contains some trace clasts with unaltered silicates (Brunetto et al., 2023), and the asteroid has been suggested to have formed near to the ice giants before its implantation into the main belt (Hopp et al., 2022; Nesvorný et al., 2024). Figure 6 shows a comparison to 67P (Kurokawa et al., 2020), a presumably non-aqueously-altered object with absorptions between 3.0-3.5 attributed to ammonium-bearing salts (Poch et al., 2020).
The 2.72 features on Ceres and Himalia are similar to some non-ammoniated meteoritic phyllosilicates. The center of the 2.7 feature caused by meteoritic phyllosilicates depends on the Mg/Fe content and therefore the degree of aqueous alteration. Initially anhydrous silicates become Fe-rich serpentine with longer wavelength centers, with continued alteration increasing abundance and replacing Fe- with Mg to form antigorite and magnetite before ultimately forming clays like saponite (Howard et al., 2011). Heavily altered CI chondrites fit farther along this progression compared with typically less altered CM chondrites, though considerable variation exists in the 3 band shapes of these groups of meteorites(Takir et al., 2019). Cere’s feature has been interpreted as similar in strength and location to CI and heavily altered CM chondrites (McSween et al., 2018). Himalia’s 2.7 band has the same center as Ceres’s, suggesting both objects have phyllosilicates with similar Mg/Fe content. The band depth and shape of Himalia’s 2.7 feature, however, may be more similar to the heavily altered CMs and C2-Ungrouped meteorites such as Tagish Lake and Essebi.
Himalia’s 2.7 absorption feature shares some similarities with meteoritic materials, but the 3.05 feature has not been detected from any known meteorite. However, Himalia’s combined 3 band complex is well-matched by laboratory samples of ammoniated phyllosilicates (De Angelis et al., 2021). Figure 7 shows that samples of illite, a clay mineral, can provide a reasonable fit to Himalia’s spectrum from 2.6-2.9 microns. However, beyond 2.9 microns, illite’s narrower absorption feature diverges from Himalia’s wider band complex. When ammoniated in a laboratory setting, illite’s absorption broadens and shows the key 3.05 micron absorption minimum. This feature survives heating up to (De Angelis et al., 2021). We note that the ammonium content on Himalia is likely significantly lower than what is contained in the ammoniated-illite sample, as Himalia’s spectrum is better described by a linear combination of untreated illite with the ammoniated sample (Figure 7.
Other, more complicated mixtures that include ammoniated materials can also fit the 2.7 and 3.05 m bands. We do not consider the case of arbitrarily many absorbers (and thus arbitrary mixture complexity), but we do note that plausible matches are also provided by mixtures of non-ammoniated phyllosilicates with a separate ammonium-bearing material. The presence of such a mixture would imply that the phyllosilicates did not interact with NH-solutions during the parent body’s aqueous processing, a geochemically incoherent situation if both materials were endogenic to the parent body. Yet this two-component example does provide a noteworthy spectral match, as illustrated in Figure 7. A combination of the hydrated C2-Ungrouped chondrite Essebi and an anhydrous ammonium-bearing material (represented by 67P’s nucleus Kurokawa et al., 2020) provides a strong fit to the shape and location of Himalia’s 3.05 feature. The Essebi/67P combination also reproduces the overall width and location of Himalia’s 3 complex, but with two caveats: Essebi contains a strong organics feature at 3.4 and has a longer wavelength band minimum.
We conclude that hydrated silicates and ammonia-bearing materials (either as ammoniated phyllosilicates or as mixed components) provides the best fit to Himalia’s band complex. We do not find any other single material that fits the overall shape of Himalia’s entire band complex from 2.7-3.8 microns. For its simplicity as a single-absorber fit, we favor ammoniated phyllosilicates as the cause of the band complex. While the laboratory samples of illite provide a useful demonstration of the ammoniation process, meteorites display a variety of phyllosilicates in their matrices. Himalia’s surface is likely to be at least similarly complex.
5 The Himalia Family’s Compositional Heterogeneity
Himalia and its family dominate the mass of the irregular satellite population, and therefore the overall abundance of materials in the system. The discovery of spectral heterogeneity within this family is surprising, particularly given the similarities between Himalia and Elara in the VNIR (Bhatt et al., 2017).
A comparison of the Himalia family members is shown in Figure 8. Himalia, Elara, and Lysithea all have major spectral features that clearly differ from each other. These differences are most easily seen by comparing Himalia and Lysithea, which have different shapes from 2.6-2.9 . Additional differences can be seen near 3 : Himalia’s rounded band is qualitatively different from Lysithea’s sharper, more jagged features in this region. Despite their distinctions, Elara suggests that these two materials are somehow linked. Elara’s absorption feature is directly matched by a simple (50/50) mean of Himalia and Lysithea’s band complexes (Figure 8, Bottom). The Himalia/Lysithea blend describes both Elara’s 2.7 feature and Elara’s intermediate band shape/center from 2.9-3.2 . Elara’s surface composition is, therefore, apparently related to both the larger Himalia and the smaller Lysithea.
One hypothesis is that Lysithea represents material that was exogenic to Himalia’s parent body, possibly sourced from the impactor of the family forming event itself. In that scenario, Elara’s intermediate spectrum would be a mixture of the impactor and the Himalian parent. However, that scenario is inconsistent with Himalia and Elara’s strong spectral similarity at all other wavelengths. Lysithea has a different color from 0.7-1.5 compared with both Elara (Figure 3) and Himalia (Bhatt et al., 2017). If bulk Lysithean material was simply combined with Himalian material, all of Elara’s spectral properties would be intermediate between Himalia and Lysithea. Instead, Elara’s 3 absorber is intermediate between Himalia and Lysithea, but its color and absorption is similar to Himalia. Furthermore, all three objects have the 3.9 feature (which could relate to carbonates). The Himalia family members, therefore, appear compositionally linked and have aqueous alteration products that vary with size.
To explain Elara’s combination of Himalia-like colors and blend of Himalian/Lysithean 3 bands, we hypothesize that these objects were all primarily formed from the same parent body. If Himalia, Elara, and Lysithea all reflect material from the same parent body, our observed progression of hydration features agrees with the hypothesis of Vilas & Hendrix (2024), who proposed that these bodies represent the parent body’s core, mantle, and crust material, respectively.
Future searches for surface heterogeneity and measurements of smaller family members are necessary to constrain family variability: it is yet unknown if there are any consistent spectral types within the Himalia family or a continuum of spectra between two endmembers. Constraining compositional trends would be useful to determine how aqueous alteration occurred on the Himalia parent body. Alteration may have been a function of depth on some irregular satellite parent bodies (e.g., Vilas & Hendrix, 2024; Sharkey et al., 2023) or perhaps occurred in more localized pockets (e.g., like the CV3 Mokoia parent body, Tomeoka & Ohnishi, 2011).
Overall, we propose that the Himalia family parent body was similar to Ceres in terms of containing water, organics, and NH-bearing materials. Furthermore, the location of Himalia’s 2.7 feature is in close agreement with the same feature on Ceres, suggesting both objects experienced similar levels of aqueous alteration. Yet Himalia and Ceres differ in the shape and depth of the 3.05 region. It is possible that the spectral distinctions reflect differences in initial abundances of NH-compounds, but Ceres and Himalia have also likely experienced divergent geological processing. Discerning whether or not Himalia and Ceres could have formed from the same primordial reservoirs provides novel opportunities to explore both objects’ migratory histories and also positions the smaller Himalia family members as possible samples of the interiors of Ceres-like objects.
The ultimate origins of Himalia, and by extension NH in the Jovian system, are still unclear. Beyond the comparisons to Ceres and other inner-solar-system asteroids, we also note similarities between Himalia and the Jovian inner satellite Amalthea from (Takato et al., 2004). Amalthea’s reflectance spectrum has not been characterized between , limiting the constraining power of this comparison considerably. However, the similarities in possible shape and location of Amalthea’s absorption band to Himalia’s are intriguing. Amalthea is thought to represent material that accreted close to Jupiter during the planet’s formation (Greenberg, 2010), and provides a critical constraint on the types of planetesimals that accreted prior to Jupiter’s hypothesized migration. Understanding if Himalia includes any Amalthea-like materials would inform whether or not Himalia itself formed near to Jupiter prior to its capture. Given the potential implications for the Himalia’s provenance, as well as the possible effects of Himalian dust on Gailelan surfaces, future spectroscopic assessment of Amalthea and the other small inner satellites is crucial to understand the Jovian satellite system as a whole.
6 Implications for Nitrogen Delivery to the Galilean moons
NH4-bearing species on Himalia and Elara represent an important source of nitrogen in the Jovian system, which could be delivered to the Galilean moons in dust grains migrating inward on slowly decaying orbits due to Poynting-Robertson drag (Burns et al., 1979). Prior ground and space-based observations have suggested that Callisto’s 4.57 feature may result from CN-bearing (or CS-bearing) compounds delivered in dust from the irregular satellites (Cartwright et al., 2020, 2024).
However, we detect no evidence for a 4.6 absorption feature, nor any reliable evidence for S-bearing species, in the irregular satellite spectra reported here. Instead, NH4-bearing compounds delivered in dust grains may provide an important source of nitrogen that mixes with H2O ice and native (or delivered) C-rich material in Callisto’s regolith. This mixture of N, C, H, and O is then irradiated by charged particles trapped in Jupiter’s magnetosphere (including S+ ions), forming new radiolytic products, possibly including OCN- and long-chain, refractory organic residues that express C-H and C-N stretching modes (e.g., Accolla et al., 2018).
Delivery and subsequent irradiation of NH4-bearing material could contribute to the subtle absorption features detected between 3.2 and 3.7 and Callisto’s broad 4.57 band. Recent dynamical modeling of dust grain trajectories indicates that dust from Jupiter’s retrograde irregular satellites should primarily strike Callisto’s leading hemisphere, whereas dust from prograde Himalia might collide with all four of the Galilean moons (Figure 8 in Chen et al. 2024). Although NH4-rich dust from Himalia could remain on the surface of Callisto and darker regions of Ganymede over long geologic timescales, this NH4-rich content might be incorporated into Europa’s interior via available geologic conduits, such as chaos-dominated terrains in Tara Regio and elsewhere, over much shorter timescales (e.g., Hand et al., 2009; Hesse et al., 2022), potentially providing an exogenic source of nitrogen to Europa’s subsurface ocean (Cartwright et al., 2024), a key element for life as we understand it.
7 Trojan-like Surfaces Exist Amongst Fragmented Irregular Satellites
Previous compositional assessment of the spectrally “red” irregular satellites have primarily relied on analogies to Jovian Trojan asteroids without specific mineralogical constraints. VNIR measurements (Sharkey et al., 2023; Grav & Holman, 2004) showed that several irregular satellites closely match Jovian Trojan color taxa. Figure 9 demonstrates color and absorption band similarities between the irregular satellites Sinope and Carme and the Jovian Trojans Leucus, Patroclus, and Eurybates (Wong et al., 2024). Sinope and Leucus are of comparable sizes (D 40 km) and have very similar spectra. Carme is an order of magnitude smaller than Patroclus. Carme and Patroclus have similar colors longward of 1.6 , but VNIR data discriminates Carme into the “red” taxon and differentiates it from the “less-red” Patroclus (Sharkey et al., 2019). This similarity in bands suggests that less-red materials could be present on redder surfaces.
The red objects (Carme, Sinope, Themisto, Figure 9, left) display similar 3.4 organics features with band depths of , consistent with Wong et al. (2024)’s suggested correlation between color and organic content. The red objects (Carme, Sinope, Themisto) have similar organic band depths of , while Patroclus’s band depth is . Carme may be consistent with less-red material that includes higher reddening from organics. Carme’s band properties suggest that it could be a transition object between red and less-red groups, or that it contains a mixture of both materials.
The presence of Trojan-like 3.0 and 3.4 absorption bands on small red Jovian irregular satellites places some new constraints on the features’ origins. The irregular satellites’ extensive collisional processing (Bottke et al., 2010) took place after their capture about Jupiter. This post-capture processing suggests that irregular satellites could have younger surfaces compared to similarly sized Jovian Trojans (as Trojans larger than are unlikely to have been substantially collisionally evolved since their own capture, Nesvorný 2018; Marschall et al. 2022; Bottke et al. 2023). In order for the same 3.0 and 3.4 absorbers to be present across a variety of irregular satellite sizes (), these materials must have been abundant enough in irregular satellite parent bodies to produce red collisional fragments (for example, if red objects have red interiors). Alternatively, red materials could be produced on satellite surfaces faster than their surfaces are collisionally refreshed. If red surfaces are formed in-place, then their formation must be related to materials that are stable within the Jovian thermal environment over those weathering timescales.
The less-red objects Pasiphae and Lysithea have 3 features which are distinct and shifted to shorter wavelengths (near ) compared to the redder objects Leucus and Sinope (bands near ). Fig. 10 shows that the band shapes observed on Lysithea and Pasiphae are good matches to those observed on Eurybates and Patroclus (Wong et al., 2024). The 3 band depths of Patroclus, Eurybates, and Lysithea/Pasiphae do not display a trend as a function of their VNIR colors.
The similarities between the absorptions on Eurybates and Pasiphae/Lysithea is particularly striking given Eurybates’ uniqueness amongst Trojans. The Eurybates collisional family is more spectrally neutral than typical Trojans and is similar in spectral slope to C-type asteroids (Roig et al., 2008). As a major outlier, it has been suggested that Eurybates may demonstrate what a younger, non-weathered Trojan surface may look like. The irregular satellites demonstrate that satellites of the same size in the same family can express a wide range of Trojan-like spectra. Figure 10 also compares the spectra of Lysithea and Eurybates to the dark, ice-poor regions of Phoebe as observed by the Cassini mission (Clark et al., 2005). This connection suggests that captured bodies at Jupiter and Saturn have similar non-ice components.
The origins of the absorber(s) among the less-red objects are also unclear. (Wong et al., 2024) noted that the 3 bands of Trojans are similar to spectra of radiolytically processed of C- and N-bearing ices. Such a scenario would be consistent with the production of 3 absorbers from irradiation in the outer solar system prior to Jovian capture, and could explain the absorptions in both red and less-red sufaces. We note an additional possibility, motivated by the discovery of Lysithea’s less-red 3 band shape within the otherwise hydrated Himalia family. Cronstedtite, the Fe-rich serpentine endmember, is the first phyllosilicate to form during aqueous alteration processes (see discussion of alteration progression by McSween et al., 2018). Terrestrial samples show an absorption band centered near to 2.8-2.9 microns (Bates et al., 2020; Gomez Barrientos et al., 2024), in potential agreement to the location of the less-red features. If any irregular satellites contained cronstedtite, that would imply they experienced a minimal but measurable amount of aqueous alteration. These materials are not represented in the meteoritic inventory, but could be analogous to petrographic grade 3 samples. We note that our comparison to Cronstedtite is based on laboratory measurements at ambient conditions (Bates et al., 2020), and contamination from adsorbed water and/or Fe-oxyhydroxides may be present that distort this absorption feature.
8 Ananke: Aqueous Alteration Among Retrograde Satellites?
While Himalia and Elara clearly demonstrate hydrated mineralogies, their bands do not resemble any known meteorites. Similarly, we do not find any clear meteorite analogs for Lysithea and Pasiphae, albeit their 3 micron bands do partially resemble some carbonaceous asteroids observed over the limited wavelength ranges accessible from previous ground-based surveys (Rivkin et al., 2015). These comparisons are severely limited by the small number of laboratory studies of relevant materials that have been prepared at vacuum to remove terrestrially adsorbed water. However, we do find several suitable comparisons relevant to Ananke.
We do not identify a specific spectral analog among the meteorite inventory, but note a broad similarity with a variety of hydrated alteration products found amongst CV, CR, and CM samples.
Fig. 11 shows that Ananke’s band is broadly consistent in shape, depth, and location with some hydrated C2 chondrites. Ananke is intermediate between CR2 Dominion Hills 10085 (Takir et al., 2019) and CM2 Murchison (Yu et al., 2024) in terms of the band center. The broad, rounded feature on Murchison is attributed to the presence of structural OH in phyllosilicates. Ananke also displays a minor absorption at 2.67 and 2.86 microns, along with a weak 3.3-3.4 micron absorption consistent with organic materials. We therefore conclude that Ananke’s 3 micron band is consistent with C2-like phyllosilicate assemblages, and that it differs strongly from the hydrated materials present on Himalia.
By suggesting that Ananke contains phyllosilicates, we provide the first evidence that aqueous alteration amongst the Jovian irregular satellites extends beyond Himalia and its prograde family. Since Ananke is in a retrograde orbital group, its parent body cannot be directly related to Himalia. Combined with Himalia, Ananke’s 3 properties suggests that parent bodies within two of the four main irregular satellite groups experienced aqueous alteration. Both the prograde and retrograde swarms include hydrated asteroids as well as red and less-red Trojan-like bodies.
9 Conclusions
We report 0.7-5.2 reflectance spectra of eight Jovian irregular satellites collected with JWST’s NIRSpec instrument. We find major variation in composition and presence of aqueous alteration products between each major orbital group. Irregular satellite families, which are presumed to have formed via collisions, contain various Trojan-like and C-type-asteroid-like surfaces throughout both prograde and retrograde populations.
The two largest irregular satellites, Himalia and Elara ( 140 km and 80 km, respectively, both prograde, Himalia group) have bands at 2.7 and 3.05 demonstrating the presence of ammoniated phyllosilicates. Elara’s major band properties are consistent with a simple mean between Himalia and Lysithea. Himalia and Elara also have a minor feature at 4.25 , similar to the feature discovered on the Trojan Eurybates and associated with complexed CO2 (Wong et al., 2024).
The observed spectral variation in the Himalia group, including Elara’s blend of Himalia-like and Lysithea-like material could relate to a stratified mutual parent body. Generally, we suggest that the complex hydration states observed within the Himalia family traces back to a heterogenous parent body, and disfavor exogenic material as the cause of the variability.
We propose that the Himalia parent body formed with materials similar to Ceres-like ammonium-bearing asteroids. Determining the distribution of ammoniated small bodies outside of the main asteroid belt is critical to determine how closely linked Himalia and Ceres may be. We note Himalia is spectrally similar to ground-based observations of the inner satellite Amalthea (Takato et al., 2004), which may represent the primitive materials that accreted near Jupiter prior to its migration. Further compositional characterization of the inner satellites is critical to assess whether Jupiter captured and retained any irregular satellites from zones near to the planet’s formation region.
At least three other distinct types of carbonaceous material are ubiquitous among both prograde and retrograde irregular satellites, with observed matches to both “red” and “less red” Jovian Trojans as well as aqueously altered carbonaceous chondrites. The retrograde satellite Ananke is similar to a variety of hydrated chondrites, suggesting that aqueous alteration is present within the retrograde satellite population and is not limited to the prograde Himalia family.
The Pasiphae family contains objects similar to red and less red Trojan asteroids. Notably, Pasiphae has a 3 absorption feature that is similar to the Himalia family member Lysithea, as well as the Trojan asteroid Eurybates. The origins of this absorber are unclear, and its presence in apparently hydrated (Himalia) and anhydrous (Pasiphae) families complicates a clear association with a single surface type.
The “red” group of satellites (Carme, Sinope, Themisto) display notable differences in their band properties. Carme has a similar band to the less-red Trojan Patroclus, and Carme’s redder color may be related to its deeper 3.4 organics band. The prograde satellite Themisto and the retrograde satellite Sinope have 3.0 and 3.4 bands similar to the red Trojan Leucus. The colors of red objects appear to correlate with the depth of the 3.4 feature. We suggest that the 3 and 3.4 absorbers on “red” irregular satellite surfaces must somehow survive the collisional disruption of their parent bodies, or that such absorbers reform on fresh surfaces faster than disruption can occur.
Ultimately, the variety of surface types and hydration states that we observe suggest that Jupiter captured satellites that include less-red and red Trojans as well as aqueously altered asteroids. This includes the capture of hydrated materials in the Ananke family and the complex materials within the Himalia family itself. The origins of the irregular satellites appear varied and require multiple progenitor populations.
10 Acknowledgments
This work is based on observations made with the NASA/ESA/CSA James Webb Space Telescope. The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. These observations are associated with program #4028. The authors wish to thank James Bauer, Jessica Sunshine, Kathryn Volk, and Shannon MacKenzie for helpful discussions that improved this work. We also thank Driss Takir for providing the original meteorite spectra used for figure comparisons.